OpenStudy (anonymous):

What is the distance between the points (3, 6) and (-21, -26) in the coordinate plane?

OpenStudy (anonymous):

\[d=\sqrt{\left( x2-x1 \right)^{2}+\left( y2-y1 \right)^{2}}\]

OpenStudy (anonymous):

32 24 30 40

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

\[d=\sqrt{\left( -21-3 \right)^{2}+\left( -26-6 \right)^{2}}=\sqrt{(-24)^{2}+\left( -32 \right)^{2}}\] \[=8\sqrt{\left( -3 \right)^{2}+\left( -4 \right)^{2}}=8\sqrt{9+16}=8*5=40\]

OpenStudy (anonymous):

thanks soooo much

OpenStudy (anonymous):

yw